# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ Abstract
We introduce a game-theoretic framework examining strategic interactions between a platform and its content creators in the presence of AI-generated content. Our model’s main novelty is in capturing creators’ dual strategic decisions: The investment in content quality and their (possible) consent to share their content with the platform’s GenAI, both of which significantly impact their utility. To incentivize creators, the platform strategically allocates a portion of its GenAI-driven revenue to creators who share their content. We focus on the class of full-sharing equilibrium profiles, in which all creators willingly share their content with the platform’s GenAI system. Such equilibria are highly desirable both theoretically and practically. Our main technical contribution is formulating and efficiently solving a novel optimization problem that approximates the platform’s optimal revenue subject to inducing a full-sharing equilibrium. A key aspect of our approach is identifying conditions under which full-sharing equilibria exist and a surprising connection to the Prisoner’s Dilemma. Finally, our simulations demonstrate how revenue-allocation mechanisms affect creator utility and the platform’s revenue.
@@ 1 Introduction
Content platforms such as YouTube, Medium, and SoundCloud have become ubiquitous, hosting vast ecosystems of creators who aim to maximize their reach, engagement, and ultimately revenue. The decision-making challenges faced by these creators has been the focus of a growing body of work, examining aspects such as revenue-sharing mechanisms [9], fairness in content visibility and diversification [8, 35], and the economic incentives shaping content creation [6, 7, 17, 2, 27]. As these platforms evolve, especially with advances in Generative AI (GenAI), new considerations arise.
@@
To address these two challenges, platforms may offer a remedy: Sharing GenAI-driven revenue with creators who allow the platform’s GenAI to use their content. This mirrors existing advertising revenue-sharing programs on platforms like YouTube and X (Twitter) [45, 56], and could incentivize both sharing consent and content quality. Therefore, the central design problem becomes: How should the platform allocate its GenAI-driven revenue to maximize its (retained) revenue while incentivizing high-quality content creation and content sharing?
@@
This paper pioneers the study on revenue sharing in the presence of strategic content creators and AI-generated content. We present a game-theoretic model consisting of a platform and strategic creators, in which the platform can generate content using GenAI tools and allocate its GenAI-driven revenue based on creator actions. Our framework incorporates creators’ dual strategic decisions: (i) the extent to which they share their content with the platform’s GenAI, ranging from full withholding to full sharing, and (ii) the body of content they wish to create, each incurring its own cost. The platform selects a revenue-sharing mechanism rewarding creators based on these two axes of contribution, with the goal of optimizing for content quality, overall revenue, or a tradeoff between them.
Our contribution
@@
Our contribution is twofold. First, we introduce the first game-theoretic model that captures both content production and data sharing in Generative AI ecosystems, where creators and the platform are strategic. User traffic is allocated through a Tullock contest structure [43], where creators and the platform’s GenAI system compete based on their relative content quality. Within this environment, each creator decides not only the quality of their content but also how much of it to share with the GenAI. The GenAI’s content quality is determined by the pool of shared creator content; thus, the platform should incentivize creators to produce and share content. To that end, we assume the platform can select a proportional allocation rule (see Equation 1) parametrized by $\rho\in[0,1]$, and redistribute a $\rho$ portion of its GenAI-driven revenue with the creators. The platform is also strategic, and can use $\rho$ in a way that steers creators towards revenue-maximizing outcomes.
@@
Our second contribution is technical. We study equilibria in the induced game, and focus on a subclass of equilibrium profiles we term full-sharing equilibria (FSE for abbreviation; see Definition 1). In an FSE, creators are better off by sharing their content. FSEs are desirable due to several properties that we discuss in Section 2.1. We present a comprehensive characterization of FSE in Section 3, showing that any $\rho$ induces at most one FSE. In Section 4, we study the platform’s revenue optimization problem (see PO). The platform can choose any value $\rho$ to maximize its revenue under the induced FSE, if it exists. Such bi-level optimization is notoriously NP-hard [26], and in our case, the difficulty is further exacerbated. Among other factors, finding equilibrium is non-trivial as best-response computation is non-concave, and the domain of feasible solutions does not have a closed-form. Nevertheless, we show that
@@
The work most closely related to ours is Yao et al. [52], who introduce a model of competition between human creators and a GenAI agent within a Tullock contest [20, 21, 23, 15], a celebrated and well-established economic model of competition. Our model, like that of Yao et al. [52], captures how user traffic is distributed based on the relative quality of content produced by humans and a GenAI system. The effectiveness of the GenAI system itself is derived from human-generated content. However, our paper significantly departs from Yao et al. [52]. First, we allow creators to decide whether and to what extent their content is shared with the GenAI system, making data sharing itself a strategic choice. Second, we model the platform as a strategic actor that designs a revenue-allocation mechanism to influence creator behavior. These modifications result in a doubly strategic setting, where both creators and the platform engage in interdependent optimization. From a technical perspective, the creators’ strategy spaces become more complex, and the game is not monotone anymore [37]. Furthermore, modeling the platform as a strategic entity gives rise to a novel bi-level optimization problem of revenue maximization, which has not been addressed in previous works. Consequently, our research questions are novel, and our results significantly extend the existing literature by addressing strategic interactions that Yao et al. [52] have not explored.
@@ 2 Model
We consider $n\geq 2$ human content creators competing for user traffic and denote the set of creators by $[n]=\{1,2,\dots,n\}$, with a typical creator indexed by $i\in[n]$. Each creator produces a body of content that encompasses multiple attributes: Quality, quantity, trend alignment, domain expertise or specialty, and so on. As is standard in prior work (e.g., [53, 52]), we aggregate these attributes into a single scalar, and without loss of generality, call it content quality. Formally, creator $i$ selects a quality $x_{i}\in[0,\infty)$.
For each creator $i$, creating a content $x_{i}$ incurs a production cost $c_{i}(x_{i})$. While we do not assume a particular form of the cost function, we have several structural assumptions that are standard in prior work. We assume that for every $i\in[n]$, the cost function $c_{i}$ is non-decreasing, twice differentiable, and strictly convex. Moreover, we assume that $c_{i}(0)=0$, $\lim_{x\to 0^{+}}\nicefrac{{dc_{i}(x)}}{{dx}}=0$, and $\lim_{x\to\infty}\nicefrac{{dc_{i}(x)}}{{dx}}=\infty$. These assumptions ensure that not creating content is costless, and that the marginal cost of improving the quality by one unit is initially zero but becomes unbounded.
@@
The platform operates a GenAI system, modeled as a non-strategic agent whose output quality depends on the aggregate quality of the content shared from creators, $\mathbf{x}^{\top}\mathbf{s}=\sum_{j=1}^{n}x_{j}s_{j}$. Its quality is given by $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=\alpha\cdot\mathbf{x}^{\top}\mathbf{s}$, Where $\alpha>0$ represents the efficiency of data usage. [^2] We discuss extensions of some of our results to a more general functional form of $Q_{\text{AI}}$ in Appendix F. Notably, if no data is shared (i.e., $\mathbf{s}=0$), then $Q_{\text{AI}}=0$.
@@
[…]
Given a strategy profile $(\mathbf{x},\mathbf{s})$ and the induced user traffic $T(\mathbf{x})$, users allocate their attention to human and AI-generated content according to a Tullock-like decision based on relative quality [43]. Specifically, the fraction of attention allocated to creator $i$’s content is $\frac{x_{i}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})}$, while the fraction allocated to AI-generated content is $\frac{Q_{\text{AI}}(\mathbf{x},\mathbf{s})}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})}$. In both cases, the denominator $\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})$ represents the total quality of created and generated content available on the platform.
@@
The platform may have multiple sources of revenue. However, to ease presentation, we focus on the user traffic its receives from its AI-generated content. Note that it depends on $Q_{\text{AI}}$, which in turns depends on human content. To obtain high-quality AI-generated content, the platform can allocate some of this revenue back to creators. Formally, the platform commits to an allocation rule $f:[0,\infty)^{n}\times[0,1]^{n}\to\Delta^{n+1}$. It maps every strategy profile $(\mathbf{x},\mathbf{s})$ to a vector $(f_{1}(\mathbf{x},\mathbf{s}),\ldots,f_{n}(\mathbf{x},\mathbf{s}),f_{P}(\mathbf{x},\mathbf{s}))$, where $f_{i}(\mathbf{x},\mathbf{s})$ represents the fraction of GenAI-driven revenue allocated to creator $i\in[n]$, and $f_{P}(\mathbf{x},\mathbf{s})$ is the fraction retained by the platform. Here, $\Delta^{n+1}$ denotes the $(n+1)$-dimensional simplex.
While various allocation mechanisms are possible, we focus on the class of parameterized proportional allocation rules $f_{\rho}$, formally defined below, and address other designs in Appendix F. The revenue-allocation parameter $\rho\in[0,1]$, or simply the allocation parameter, specifies the fraction of GenAI-driven revenue allocated to creators, with the remainder $1-\rho$ retained by the platform. Under this rule, the platform distributes the $\rho$-fraction among creators in proportion to their effective shared contributions $x_{i}s_{i}$, yielding
$f_{i,\rho}(\mathbf{x},\mathbf{s})=\begin{cases}\rho\cdot\frac{x_{i}s_{i}}{\mathbf{x}^{\top}\mathbf{s}}&\mathbf{x}^{\top}\mathbf{s}>0,\\ 0&\mathbf{x}^{\top}\mathbf{s}=0\end{cases}$ (1)
@@ 2.1 Solution Concepts
We are interested in stable outcomes in the subgame $G(\cdot)$, where stability refers to pure Nash equilibrium (PNE). We highlight a class of PNEs that we call Full Sharing PNE, or FSE for abbreviation. In FSEs, full content sharing arises endogenously: Each creator chooses to share their content because the platform’s revenue-allocation mechanism renders sharing a best response.
@@
To motivate Definition 2, imagine the auxiliary enforced sharing game, where $f_{\rho}$ is fixed and creators must share their content, i.e., $\mathbf{s}=\mathbf{1}$. ESEs allow us to analyze game dynamics by focusing solely on creators’ quality profiles $\mathbf{x}$. We remark that any quality profile $\mathbf{x}$ such that $(\mathbf{x},\mathbf{1})$ is an FSE also constitutes an ESE.
3 FSE Characterization
In this section, we analyze how creators respond to an allocation rule. Subsection 3.1 assumes a fixed allocation rule $f_{\rho}$ and studies the structure of ESEs in $G(\rho)$. We show that every $\rho$ induces a unique ESE, although different values of $\rho$ may induce different ESEs. In Subsection 3.2, we move to study FSEs. Namely, Theorem 5 characterizes a sufficient condition under which the unique ESE $\mathbf{x}^{\star}$ also constitutes the unique FSE $(\mathbf{x}^{\star},\mathbf{1})$.
@@ 3.1 Best Sharing Response and Uniqueness of ESE
Proposition 2.
Fix an arbitrary creator $i$ with a content $x_{i}>0$, and a profile $(\mathbf{x}_{-i},\mathbf{s}_{-i})$. Further, let $\tau_{i}$ such that $\tau_{i}=\frac{x_{i}}{x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}$. The optimal sharing level $s$ of creator $i$ w.r.t. $(x_{i},\mathbf{x}_{-i},\mathbf{s}_{-i})$, i.e., $s\in\operatorname*{arg\,max}_{s\in[0,1]}U_{i}((x_{i},\mathbf{x}_{-i}),(s,\mathbf{s}_{-i});f_{\rho})$ is given by $s=\begin{cases}1&\rho>\tau_{i}\\ \textnormal{any number from }[0,1]&\rho=\tau_{i}\\ 0&\rho<\tau_{i}\\ \end{cases}$.
Proposition 2 reveals an interesting asymmetry: Creators producing a large share of $\left\|\mathbf{x}\right\|$ have higher thresholds; thus, they are more inclined to withhold their data, as doing so weakens a powerful GenAI competitor. In contrast, creators with relatively low quality have stronger incentives to share. Notably, the sharing decision becomes effectively binary—each creator either fully shares or fully withholds, with no incentive to choose an intermediate level.
Another insight from Proposition 2 concerns the case $\mathbf{s}=\mathbf{0}$, where all creators initially withhold their data. Fix a quality profile $\mathbf{x}$, and notice that if the number of creators $n$ is moderately large, the threshold $\tau_{i}=\frac{x_{i}}{\left\|\mathbf{x}\right\|}$ becomes small for most creators. Consequently, even a low revenue-allocation parameter $\rho\approx\nicefrac{{1}}{{n}}$ can make it beneficial for the least productive creator, denoted $i$, to deviate and set $s_{i}=1$. This deviation increases $\mathbf{x}^{\top}\mathbf{s}$, thereby lowering the threshold $\tau_{j}$ for other creators $j\neq i$ to $\frac{x_{j}}{\left\|\mathbf{x}\right\|+\alpha x_{i}}<\frac{x_{j}}{\left\|\mathbf{x}\right\|}=\tau_{j}$. As a result, creator $i$’s deviation can potentially trigger a cascade of full-sharing responses from other players, an effect similar to the classic prisoner’s dilemma.
Next, we examine creators’ quality choices under enforced sharing.
Proposition 3.

### v2

# Strategic Content Creation with GenAI: To Share or Not to Share?

@@ Abstract
We introduce a game-theoretic framework that examines strategic interactions between a platform and its content creators in the presence of AI-generated content. Our model’s main novelty is in capturing creators’ dual strategic decisions: their investment in content quality and their (possible) consent to share their content with the platform’s GenAI, both of which significantly impact their utility. To incentivize creators, the platform strategically allocates a portion of its GenAI-driven revenue to creators who share their content. We focus on the class of full-sharing equilibrium profiles, in which all creators willingly share their content with the platform’s GenAI system. Such equilibria are highly desirable both theoretically and practically. Our main technical contribution is formulating and efficiently solving a novel optimization problem that approximates the platform’s optimal revenue subject to inducing a full-sharing equilibrium. A key aspect of our approach is identifying conditions under which full-sharing equilibria exist and a surprising connection to the Prisoner’s Dilemma. Finally, our simulations demonstrate how revenue-allocation mechanisms affect creator utility and the platform’s revenue.
@@ 1 Introduction
Online Content platforms such as YouTube, Medium, and SoundCloud have become ubiquitous, hosting vast ecosystems of creators who aim to maximize their reach, engagement, and revenue. The decision-making challenges faced by these creators have been the focus of a growing body of work, examining aspects such as revenue-sharing mechanisms [9], fairness in content visibility and diversification [8, 33], and the economic incentives shaping content creation [7, 6, 15, 2, 25]. As these platforms evolve, especially with advances in Generative AI (GenAI), new considerations arise.
@@
To address these challenges, platforms may offer a remedy: sharing GenAI-driven revenue with creators who allow the platform’s GenAI to use their content. This mirrors existing revenue-sharing programs on platforms like YouTube and X (formerly Twitter) [44, 55], and could incentivize both sharing consent and content quality. The main question thus becomes: how should the platform allocate its GenAI-driven revenue to maximize its (retained) revenue while incentivizing the creation and sharing of high-quality content?
@@
This paper pioneers the study of revenue sharing in the presence of strategic content creators and AI-generated content. We present a game-theoretic model consisting of a platform and strategic creators, in which the platform can generate content using GenAI tools and allocate its GenAI-driven revenue based on creator actions. Our framework incorporates creators’ dual strategic decisions: (i) the extent to which they share their content with the platform’s GenAI, ranging from full withholding to full sharing, and (ii) the body of content they create, each incurring its own cost. The platform selects a revenue-sharing mechanism rewarding creators based on these two axes of contribution, with the goal of optimizing for content quality, overall revenue, or a tradeoff between them.
1.1 Our Contribution
@@
Our contribution is twofold. First and foremost, we introduce the first game-theoretic model that captures both content production and data sharing in Generative AI ecosystems, where creators and the platform are strategic. User traffic is allocated through a Tullock contest structure [42], where creators and the platform’s GenAI system compete based on their relative content quality. Within this environment, each creator decides not only the quality of their content but also how much of it to share with the GenAI. The GenAI’s content quality is determined by the pool of shared creator content; thus, the platform should incentivize creators to produce and share content. To that end, we assume the platform can select a proportional allocation rule (see Equation˜1) parametrized by $\rho\in[0,1]$, and redistribute a $\rho$-portion of its GenAI-driven revenue to the creators. The platform is also strategic, and can use $\rho$ in a way that steers creators towards revenue-maximizing outcomes.
@@
Our second contribution is technical. We study equilibria in the induced game, and focus on a subclass of equilibrium profiles we term full-sharing equilibria (FSE for abbreviation; see Definition 1). In an FSE, creators are better off by sharing their content. FSEs are desirable due to several properties that we discuss in Section˜2.1. We present a comprehensive characterization of FSE in Section˜3, showing that any $\rho$ induces at most one FSE. In Section˜4, we study the platform’s revenue optimization problem (see (PO)). The platform can choose any value $\rho$ to maximize its revenue under the induced FSE, if it exists. Such bilevel optimization is notoriously NP-hard [24], and in our case, the difficulty is further exacerbated. Among other factors, finding equilibrium is non-trivial as best-response computation is non-concave, and the domain of feasible solutions does not have a closed-form. Nevertheless, we show the following:
@@ 1.2 Related Work
The work most closely related to ours is Yao et al. [50], who introduce a model of competition between human creators and a GenAI agent within a Tullock contest [19, 20, 22, 14], a celebrated and well-established economic model of competition. Our model, like that of Yao et al. [50], captures how user traffic is distributed based on the relative quality of content produced by humans and a GenAI system. The effectiveness of the GenAI system itself is derived from human-generated content. However, our paper significantly departs from Yao et al. [50]. First, we allow creators to decide whether and to what extent their content is shared with the GenAI system, making data sharing itself a strategic choice. Second, we model the platform as a strategic actor that designs a revenue-allocation mechanism to influence creator behavior. These modifications result in a doubly strategic setting, where both creators and the platform engage in interdependent optimization. From a technical perspective, the creators’ strategy spaces become more complex, and the game is no longer monotone [35]. Furthermore, modeling the platform as a strategic entity gives rise to a novel bilevel optimization problem of revenue maximization, which has not been addressed in previous works. Consequently, our research questions are novel, and our results significantly extend the existing literature by addressing strategic interactions that Yao et al. [50] have not explored.
@@ 2 Model
We consider $n\geq 2$ human content creators competing for user traffic and denote the set of creators by $n]=\{1,2,\dots,n\}$, with a typical creator indexed by $i\in[n]$. Each creator produces a body of content that encompasses multiple attributes: quality, quantity, trend alignment, domain expertise or specialty, and so on. As is standard in prior work (e.g., [52, 50]), we aggregate these attributes into a single scalar, and without loss of generality, call it content quality. Formally, creator $i$ selects a quality $x_{i}\in[0,\infty)$.
For each creator $i$, creating a content $x_{i}$ incurs a production cost $c_{i}(x_{i})$. While we do not assume a particular form of the cost function, we have several structural assumptions that are standard in prior work. We assume that for every $i\in[n]$, the cost function $c_{i}$ is non-decreasing, twice differentiable, and strictly convex. Moreover, we assume that $c_{i}(0)=0$, $\lim_{x\to 0^{+}}\frac{dc_{i}(x)}{dx}=0$, and $\lim_{x\to\infty}\frac{dc_{i}(x)}{dx}=\infty$. These assumptions ensure that not creating content is costless, and that the marginal cost of improving the quality by one unit is initially zero but becomes unbounded.
@@
The platform operates a GenAI system, modeled as a non-strategic agent whose output quality depends on the aggregate quality of the content shared from creators, $\mathbf{x}^{\top}\mathbf{s}=\sum_{j=1}^{n}x_{j}s_{j}$. Its quality is given by $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=\alpha\cdot\mathbf{x}^{\top}\mathbf{s}$, where $\alpha>0$ represents the efficiency of data usage. [^2] We discuss extensions of some of our results to a more general functional form of $Q_{\text{AI}}$ in Section 6. Notably, if no data is shared (i.e., $\mathbf{s}=0$), then $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=0$. Whenever $\mathbf{x}$ and $\mathbf{s}$ are clear from the context, we simplify the notation and use $Q_{\text{AI}}$.
@@
The platform attracts user traffic denoted by $T(\mathbf{x})$, which increases monotonically with the total quality of human content, $\left\|\mathbf{x}\right\|=\sum_{j=1}^{n}x_{j}$, where we use $\|\cdot\|$ without a subscript to denote the $\ell_{1}$ norm throughout the paper. We model traffic as $T(\mathbf{x})=\mu\cdot\left\|\mathbf{x}\right\|^{\gamma}$, where $\mu>0$ is a baseline traffic parameter and $\gamma\in[0,1]$ captures the elasticity of traffic with respect to the total human content quality [39, 13]. [^3] We assume that users are attracted to the platform due to its human-generated content, not the AI-generated content. The reason is that users tend to value originality and exhibit bias against AI-generated content when it lacks it [5], which is the case in emerging domains where historical data is scarce [32]. We discuss extensions of our results to other form of traffic in Section 6.
Given a strategy profile $(\mathbf{x},\mathbf{s})$ and the induced user traffic $T(\mathbf{x})$, users allocate their attention to human and AI-generated content according to a Tullock-like decision based on relative quality [42]. Specifically, the fraction of attention allocated to creator $i$’s content is $\frac{x_{i}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})}$, while the fraction allocated to AI-generated content is $\frac{Q_{\text{AI}}(\mathbf{x},\mathbf{s})}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})}$. In both cases, the denominator represents the total quality of created and generated content available on the platform.
@@
The platform may have multiple sources of revenue. However, to ease presentation, we focus on the user traffic it receives from its AI-generated content. Note that it depends on $Q_{\text{AI}}$, which in turn depends on human content. To obtain high-quality AI-generated content, the platform can allocate some of this revenue back to creators. Formally, the platform commits to an allocation rule $f:[0,\infty)^{n}\times[0,1]^{n}\to\Delta^{n+1}$. It maps every strategy profile $(\mathbf{x},\mathbf{s})$ to a vector $(f_{1}(\mathbf{x},\mathbf{s}),\ldots,f_{n}(\mathbf{x},\mathbf{s}),f_{P}(\mathbf{x},\mathbf{s}))$, where $f_{i}(\mathbf{x},\mathbf{s})$ represents the fraction of GenAI-driven revenue allocated to creator $i\in[n]$, and $f_{P}(\mathbf{x},\mathbf{s})$ is the fraction retained by the platform. Here, $\Delta^{n+1}$ denotes the $(n+1)$-dimensional simplex.
While various allocation mechanisms are possible, we focus on the class of parameterized proportional allocation rules $f_{\rho}$, formally defined below, and address other designs in Appendix˜E. The revenue-allocation parameter $\rho\in[0,1]$, or simply the allocation parameter, specifies the fraction of GenAI-driven revenue allocated to creators, with the remainder $1-\rho$ retained by the platform. Under this rule, the platform distributes the $\rho$-fraction among creators in proportion to their effective shared contributions $x_{i}s_{i}$, yielding
$f_{i,\rho}(\mathbf{x},\mathbf{s})=\begin{cases}\rho\cdot\frac{x_{i}s_{i}}{\mathbf{x}^{\top}\mathbf{s}}&\mathbf{x}^{\top}\mathbf{s}>0\\ 0&\mathbf{x}^{\top}\mathbf{s}=0\end{cases},$ (1)
@@ 2.1 Solution Concepts
We are interested in stable outcomes in the subgame $G(\cdot)$, where stability refers to pure Nash equilibrium (PNE). We highlight a class of PNEs that we call Full Sharing PNE, or FSE for abbreviation. In FSE, full content sharing arises endogenously: each creator chooses to share their content because the platform’s revenue-allocation mechanism renders sharing a best response.
@@
To motivate Definition˜2, imagine the auxiliary enforced sharing game, where $f_{\rho}$ is fixed and creators must share their content, i.e., $\mathbf{s}=\mathbf{1}$. ESEs allow us to analyze game dynamics by focusing solely on creators’ quality profiles $\mathbf{x}$. We remark that any quality profile $\mathbf{x}$ such that $(\mathbf{x},\mathbf{1})$ is an FSE also constitutes an ESE.
3 FSE Analysis
In this section, we analyze how creators respond to an allocation rule. Subsection 3.1 assumes a fixed allocation rule $f_{\rho}$ and studies the structure of ESEs in $G(\rho)$. We show that every $\rho$ induces a unique ESE, although different values of $\rho$ may induce different ESEs. In Subsection 3.2, we move to study FSEs. Namely, Theorem˜5 characterizes a sufficient condition under which the unique ESE $\mathbf{x}^{\star}$ also constitutes the unique FSE $(\mathbf{x}^{\star},\mathbf{1})$.
@@ 3.1 Best Sharing Response and Uniqueness of ESE
Lemma 2.
Fix an arbitrary creator $i$ with content $x_{i}>0$, and a profile $(\mathbf{x}_{-i},\mathbf{s}_{-i})$. Further, let $\tau_{i}$ such that $\tau_{i}=\frac{x_{i}}{x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}}$. The optimal sharing level $s$ of creator $i$ w.r.t. $(x_{i},\mathbf{x}_{-i},\mathbf{s}_{-i})$, i.e., $s\in\operatorname*{arg\,max}_{s\in[0,1]}U_{i}((x_{i},\mathbf{x}_{-i}),(s,\mathbf{s}_{-i});f_{\rho})$ is given by
$s^{\star}=\begin{cases}1&\rho>\tau_{i}\\ \textnormal{any number from }[0,1]&\rho=\tau_{i}\\ 0&\rho<\tau_{i}\\ \end{cases}.$ (2)
Due to space constraints, all proofs in this and subsequent sections are deferred to the appendix. The proof of Lemma˜2 involves showing that creator $i$’s utility is strictly increasing in $s_{i}$ when $\rho>\tau_{i}$, strictly decreasing when $\rho<\tau_{i}$, and constant when $\rho=\tau_{i}$, yielding Equation˜2.
Lemma˜2 reveals an interesting asymmetry: creators producing a large share of $\left\|\mathbf{x}\right\|$ have higher thresholds; thus, they are more inclined to withhold their data, as doing so weakens a powerful GenAI competitor. In contrast, creators with relatively low quality have stronger incentives to share. Notably, the sharing decision becomes effectively binary—each creator either fully shares or fully withholds, with no incentive to choose an intermediate level.
Another insight from Lemma˜2 concerns the case $\mathbf{s}=\mathbf{0}$, where all creators initially withhold their data. Fix a quality profile $\mathbf{x}$, and notice that if the number of creators $n$ is moderately large, the threshold $\tau_{i}=\frac{x_{i}}{\left\|\mathbf{x}\right\|}$ becomes small for most creators. Consequently, even a low revenue-allocation parameter $\rho\approx\nicefrac{{1}}{{n}}$ can make it beneficial for the least productive creator, denoted by $i$, to deviate and set $s_{i}=1$. This deviation increases $\mathbf{x}^{\top}\mathbf{s}$, thereby lowering the threshold $\tau_{j}$ for other creators $j\neq i$ to $\frac{x_{j}}{\left\|\mathbf{x}\right\|+\alpha x_{i}}<\frac{x_{j}}{\left\|\mathbf{x}\right\|}=\tau_{j}$. As a result, creator $i$’s deviation can potentially trigger a cascade of full-sharing responses from other players, an effect similar to the classic prisoner’s dilemma.
Next, we study creators’ quality choices under enforced sharing.
Lemma 3.
@@
Given the existence and uniqueness established in Lemma˜3, we can define the ESE correspondence: let $\mathcal{X}:[0,1]\to[0,\infty)^{n}$ denote the mapping from $\rho$ to its corresponding unique ESE profile, $\mathcal{X}(\rho)$, in the subgame $G(\rho)$.